OpenStudy (anonymous):

Solve the equation x+ 1/6 = 6 a. x = -6 1/6 b. x = -5 5/6 c. x =5 5/6 d. x =6 1/6

OpenStudy (maddy1251):

When solving equations like this, you want 'x' to be alone, because that is what we are solving for.

OpenStudy (maddy1251):

\[x+\frac{ 1 }{ 6 } = 6\]

OpenStudy (maddy1251):

In order to get 'x' alone, we will subtract 1/6 from each side, because we are performing the opposite of what 1/6 is doing to that side of the equation.

OpenStudy (maddy1251):

\[x+\frac{ 1 }{ 6 } - \frac{ 1 }{ 6 } = 6 - \frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

okay thank u I gave you a medal I get it thank u

OpenStudy (maddy1251):

Thank you. :-) Make sure you perform out all operations in math equations to get a final answer.

OpenStudy (anonymous):

so is it a

OpenStudy (maddy1251):

No, C.\[6 - \frac{ 1 }{ 6 } = 5 \frac{ 5 }{ 6 }\]

OpenStudy (maddy1251):

You want to finish out the operations.